OpenStudy (anonymous):

rationalize 1/(sqrt x)

OpenStudy (unklerhaukus):

\[\frac 1{\sqrt x}=\frac 1{\sqrt x}\times1=\frac 1{\sqrt x}\times\frac {\sqrt x}{\sqrt x}=\]